solve the following radicals
sq rt x-4= sq rt 2x-5

Answer by checkley77(12844)

sqrt(x-4)=sqrt(2x-5)
squaring both sides we get:
x-4=2x-5
x-2x=-5+4
-x=-1
x=1 answer.,
proof:
sqrt(1-4)=sqrt(2*1-5)
sqrt(-3)=sqrt(2-5)
sqrt(-3)=sqrt(-3)